# A Blue Start: A large-scale pairwise and higher-order social network dataset

### v1

@@
Abstract
@@
Large-scale networks have been instrumental in shaping the way that we think about how individuals interact with one another, developing key insights in mathematical epidemiology, computational social science, and biology. However, many of the underlying social systems through which diseases spread, information disseminates, and individuals interact are inherently mediated through groups of arbitrary size, known as higher-order interactions. There is a gap between higher-order dynamics of group formation and fragmentation, contagion spread, and social influence and the data necessary to validate these higher-order mechanisms. Similarly, few datasets bridge the gap between these pairwise and higher-order network data. Because of its open API, the Bluesky social media platform provides a laboratory for observing social ties at scale. In addition to pairwise following relationships, unlike many other social networks, Bluesky features user-curated lists known as ”starter packs” as a mechanism for social network growth. We introduce ”A Blue Start”, a large-scale network dataset comprising 26.7M users and their 1.6B pairwise following relationships and 301.3K groups representing starter packs. This dataset will be an essential resource for the study of higher-order network science.
I Background & Summary
@@
Network science has revolutionized the study of human behavior because of its ability to represent complex relationships between individuals in a population. Representing social structures as networks offers insights on the temporal nature of interactions [1], the heterogeneity of interaction patterns [2, 3, 4], preferential relationships between nodes with similar characteristics [5, 6] and highly connected nodes [7], and community structure [8]. One particular medium that spawned an entire sub-field of network science was Twitter, the social media platform. This online ”laboratory” was particularly useful because of its open automated programming interface (API); the free ”Twitter Streaming API” provided researchers with up to 1% of all tweets, while the paid ”Twitter Firehose” allowed researchers to extract 100% of all public posts [9]. As Twitter grew in size, the Firehose API was discontinued and replaced with the Decahose API which allowed researchers to download 10% of all posts in real time [10, 11]. Nevertheless, the relative availability of Twitter data, along with the platform’s popularity and use across demographic and sociographic groups [12] spurred scientific innovations across a number of disciplines, including epidemic modeling [13], political communication [14], and statistical physics [15], to name just a few. Some sub-disciplines that depend on access to ”everyday” communication, such as social movement studies, were revolutionized by access to the Twitter API, which provided data to support hundreds of papers on the structure and discursive content of social movements around the world (e.g., see Refs. [16, 17, 18]).
@@
The change in ownership of Twitter, subsequently known as X, spawned interest in many alternative platforms such as Mastodon, Threads, and Bluesky [19]. We focus our attention of Bluesky because of its inherent open access to platform data. Bluesky originated in 2019 as a research initiative at Twitter examining the potential for open-sourcing and decentralizing social media [20]. It was incorporated as an independent company in 2021  [21]. Bluesky’s central innovation is a standard known as the Authenticated Transfer (AT) Protocol [22]. The AT Protocol emphasizes federated identities, allowing users to own and curate a single online identity that is usable on all AT Protocol platforms and services [23]. User data is stored in repositories on personal data servers (PDSs) that log user actions; these repositories are intended to be portable so that users can easily migrate between hosts [24]. The Bluesky application queries a user’s PDS to deliver content to users; however, anyone can build an application to view AT Protocol data. Applications with access to the same data can vary in their content moderation policies, post visibility, or algorithmic curation strategies, for example [25].
@@
This Bluesky dataset addresses both of these gaps in the corpus of higher-order network datasets. In the summer of 2024, Bluesky developers announced a new feature in the Bluesky app, known as starter packs (SPs). Starter packs are user-curated lists of accounts and feeds (which serve post content to users based on custom algorithms) with a user-supplied title and description. Users share these lists so that other users can quickly grow the list of accounts they follow based on specified interest. One has the option to either browse the list and cherry-pick accounts to follow or hit the ”follow all” button to follow all accounts in the list. Starter packs are inherently higher-order. At the time of this writing, users can create starter packs with 8 to 150 users and feeds. There is a caveat, however, which is that a similar higher-order structure on Bluesky known as ”lists”, which are not size restricted and so contain anywhere between 0 and several thousand items. With the advent of starter packs, users developed tools for converting lists into starter packs, bypassing the standard size limits. So while the list of starter packs is almost entirely higher-order, there are exceptions to this (as can be later seen in Fig. 1). One can also extract the list of accounts that each user follows, which can be represented as a directed, pairwise network. These two types of relationships bridge the gap between pairwise and higher-order network data. Furthermore, this dataset offers researchers ground to investigate social phenomena exclusive to groups phenomena and develop measures and algorithms tailored to such structures. In particular, the absence of edges in the following network when compared to the network induced by starter packs facilitates exploration of patterns that are exclusive to higher-order networks. For example, one can analyze user overlap patterns between starter packs, a structural feature inaccessible to pairwise network analysis. Higher-order network analysis enables a deeper understanding of the complex dynamics among users participating in group interactions.
Beyond applications for the study of higher-order dynamics and interactions, our data enables novel research in communication and human-computer interaction. Future work in this area can leverage our dataset to address questions of structural differences in start pack inclusion, the effect of starter pack inclusion on following relationships, and more. These questions also lend themselves to salient new possibilities in graph machine learning. In this setting, empirical data on starter packs becomes crucial for training and testing models. More generally, the broad applicability of online social network data across scientific fields is well documented in the extensive literature that has emerged over the past decade, leveraging existing datasets from other platforms. Now, with up-to-date data becoming increasingly difficult to access, this dataset provides a unique opportunity to gain early insights into a key, distinguishing feature of a burgeoning new online social media platform.
II Methods
II.1 Obtaining the list of users
The complete list of users and all operations performed on a user’s account, such account creation, modification, and deletion, are stored in the ”DID PLC Directory” at https://plc.directory. User accounts are assigned a ”decentralized identifier” (DID) which provides a persistent global reference for all account information and all queries to the Bluesky API to collect user information such as the accounts followed or the list of starter packs require a DID as a query parameter. Although in principle one could start from a given user and their associated DID and perform a breadth-first search through the following relationships to extract the list of users, the network is disconnected and, as seen in Fig. 5(b), some of the strongly-connected components are non-negligible in size. The export endpoint enables efficient downloading of the contents of the PLC Directory and we extracted the list of unique DIDs from the exported data in batches of 1,000. Although the PLC Directory also tracks account deactivation or deletion, we compiled the list of DIDs without considering this information. We collected the list of DIDs on January 22nd, 2025.
II.2 Collecting starter packs
We used the public app.bsky.graph.getActorStarterPacks public API endpoint to obtain the list of all starter packs associated with a user account. For each unique account DID extracted from the directory mentioned above, we queried this endpoint and extracted the Uniform Resource Identifier (URI) corresponding to the list representation of the starter pack for each starter pack the user created. Next, we queried the app.bsky.graph.getList public API endpoint with each URI associated with the user account and obtained the following information on each starter pack: (1) the name and description of the starter pack; (2) the creation date; (3) the starter pack creator’s DID, handle, and account creation date; and (4) the list of user accounts and feeds comprising the starter packs. Querying the app.bsky.graph.getList endpoint returns a cursor which allowed us to page through in batches of 100 starter pack items and collect the entire list of entities in the starter pack. The name and description of a starter pack are optional; if a user does not specify a name, it defaults to ”{user display name}’s Starter Pack” and if a user does not write a description, it is left empty. Note that entities in a starter pack are not necessarily users, they can be any entity that can be ”followed” on Bluesky (e.g. a list or a feed). Lists can only be followed, however, by circumventing the standard user interface through the API. We used the Python asyncio and aiohttp packages to efficiently download data from the API while respecting the rate limits [^1] https://docs.bsky.app/docs/advanced-guides/rate-limits imposed by the Bluesky API. We downloaded data on each starter pack as a Python dictionary and appended each pack to a list of all starter packs. We then saved this list to a JSON blob, described more in Section III.1, and used the XGI Python package [58] to convert the starter pack network to the HIF standard [^2] https://github.com/pszufe/HIF-standard and a hyperedge list. We collected the list of starter packs over several days ending on February 3rd, 2025.
II.3 Collecting network data
For each account DID collected from querying the DID PLC directory, we obtained the list of users followed by that account by querying the app.bsky.graph.getFollows public API endpoint. Querying this endpoint returns a cursor which allowed us to page through batches of 100 following relationships and collate the list of all accounts followed. This process yielded a snapshot of all following relationships for all accounts corresponding to the DID PLC directory. We used asynchronous programming, as explained in Section II.2, to efficiently download the data. We collected the user following relationships over several days ending on February 3rd, 2025.
II.4 Anonymizing user data
@@
Although all user data is available via the public Bluesky API and users can require authentication to access their public user data from the API, we nonetheless take reasonable steps to prevent re-identification of users. We recognize, however, that with enough effort, that it is possible to re-identify all users. However, we assert that this effort will be comparable to querying the API from scratch. We anonymize our dataset in three ways: (1) by converting the DIDs of all users and feeds to unique integer identifiers, (2) by removing the name and description of all starter packs, and (3) by coarsening the precision of all timestamps to one day. We anonymize the list of starter pack members and user accounts by first taking the union of the list of entities in the higher-order network of starter packs and the list of all user accounts, then assigning integers to every listed entity. We do the same for each starter pack URI, converting it to an integer. Then we iterate through every starter pack and every following relationship using this mapping from DIDs and URIs to unique integers. This process ensures that while users cannot be identified, node IDs are consistent across both the following network and the starter pack network.
II.5 Limitations
@@
The approach we utilized in collecting the starter pack and following data comes with limitations, however. The number of active accounts, starter packs, and following relationships on the Bluesky platform is rapidly changing and any collection method not taking this into account will not perfectly capture the network structure. Using the PLC SIS Directory as a static directory does not consider accounts created after this list was collected. This leads to collected DIDs in the starter pack and following data that are not present in the ”global” list of DIDs; existing users can both follow and add to a starter pack users that have been created since the list of DIDs was collected. Secondly, time stamps and user data can be arbitrarily set by users if they interact with Bluesky directly through the AT Protocol and not through the Bluesky application. Lastly, follows (and starter pack members) can be added and removed while waiting to process a particular DID and similarly, after that DID has been processed, following relationships and curated starter packs can change as well. One way around this is by downloading streaming data from the Bluesky Firehose API while backfilling all events for all users, as is done in Ref. [29]. The downside to this approach is that it is an immense amount of data; roughly 27GB of JSON data per day at the time of writing. We believe that our approach strikes the balance of being computationally manageable, while surveying close to the entire network.
II.6 Ethical Considerations
@@
This data collection was ruled an exempt protocol by the Institutional Review Board (IRB) at Northeastern University (No. 25-01-22) and by the Institutional Review Board for the Social and Behavioral Sciences (IRB-SBS) at the University of Virginia (No. 7332). All of the data obtained is publicly accessible (without authentication or a Bluesky account) via the Bluesky API. Additionally, this study was granted a waiver of informed consent from both the IRB at Northeastern University and the IRB-SBS at the University of Virginia; arguably, obtaining informed consent for all users’ inclusion in this dataset would constitute a larger risk to users’ privacy than the collection of publicly available data. Nonetheless, we recognize that this dataset could enable large-scale identification of accounts’ following relationships. We therefore redact all user DIDs (decentralized identifiers, used for long-term, persistent identification of entities in the AT Protocol that Bluesky uses) and truncate the timestamps on all starter pack and account creation events to include only the date; i.e., the level of precision at which timestamps are reported is one day. The details of this process are described in Section II.4.
III Data Records
This study releases two data records: (1) a list of starter packs on Bluesky and (2) a list of user following relationships.
III.1 Bluesky starter packs
@@
JSON Blob (deidentified_starterpack_blob.json, 132.7 MB). We provide a JSON file of all scraped starter packs’ contents and metadata, intended for computational analysis. This file contains a list of starter packs; each starter pack contains the fields described in Table 1.
@@
creator-date-created date The date on which the SP creator’s account was created
@@
members list of integers Each entry is a member user’s anonymous integer identifier
@@
Hyperedge list (deidentified_starterpack_edgelist.csv, 85.1 MB). We also provide a hyperedge list stored as a CSV file. Each line in this file corresponds to a starter pack, and each entry in the line is a member of the starter pack. Each line contains entries separated by commas; a single line corresponds to a single starter pack. For example, the line ”1,2,3,4” indicates that users with anonymous integer identifiers 1, 2, 3, and 4 belong to the same starter pack.
Hypergraph Interchange Format (deidentified_starterpack_hif.json, 619.1 MB). This file is a HIF-compliant JSON file. The HIF standard provides a cross-platform standard for sharing attributed higher-order network data. The edge attributes (”pack-id”, ”creator-id”, ”creator-date-created”, and ”date-created”) are labeled as described in Table 1.
III.2 Bluesky following network
The following network dataset is stored in the file deidentified_follows_edgelist.zip (8.5 GB) as a compressed CSV file. Each line $(A,B)$ in the file indicates that user $A$ follows user $B$. Here, $A$ and $B$ are anonymous integer identifiers that map to user DIDs. As described in Section II.4, these anonymous integer IDs can be linked to creators and members of starter packs.
IV Technical Validation

### v2

@@
A Blue Start: A large-scale pairwise and higher-order social network dataset
Alyssa Smith smith.alyss@northeastern.edu Network Science Institute, Northeastern University, Boston, Massachusetts, USA Ilya Amburg ilya.amburg@pnnl.gov Pacific Northwest National Laboratory, Seattle, Washington, USA Sagar Kumar kumar.sag@northeastern.edu Network Science Institute, Northeastern University, Boston, Massachusetts, USA Center for Health Informatics, Boston Children’s Hospital, Boston, Massachusetts, USA Brooke Foucault Welles b.welles@northeastern.edu Network Science Institute, Northeastern University, Boston, Massachusetts, USA Department of Communication Studies, Northeastern University, Boston, Massachusetts, USA Nicholas W. Landry nicholas.landry@virginia.edu Department of Biology, University of Virginia, Charlottesville, Virginia, USA School of Data Science, University of Virginia, Charlottesville, Virginia, USA Vermont Complex Systems Institute, University of Vermont, Burlington, Vermont, USA
@@ A Blue Start: A large-scale pairwise and higher-order social network dataset
Large-scale networks have been instrumental in shaping how we think about social systems, and have undergirded many foundational results in mathematical epidemiology, computational social science, and biology. However, many of the social systems through which diseases spread, information disseminates, and individuals interact are inherently mediated through groups, known as higher-order interactions. A gap exists between higher-order models of group formation and spreading processes and the data necessary to validate these mechanisms. Similarly, few datasets bridge the gap between pairwise and higher-order network data. The Bluesky social media platform is an ideal laboratory for observing social ties at scale through its open API. Not only does Bluesky contain pairwise following relationships, but it also contains higher-order social ties known as “starter packs” which are user-curated lists designed to promote social network growth. We introduce “A Blue Start”, a large-scale network dataset comprising 39.7M user accounts, 2.4B pairwise following relationships, and 365.8K groups representing starter packs. This dataset will be an essential resource for the study of higher-order networks.
Background & Summary
@@
Network science has revolutionized the study of human behavior through its ability to represent complex relationships between individuals in a population. Representing social structures as networks offers insights on the temporal nature of interactions [13], the heterogeneity of interaction patterns [17, 6, 25], preferential relationships between nodes with similar characteristics [50, 51] and highly connected nodes [68], and community structure [23]. One particular medium that spawned an entire sub-field of network science was Twitter, the social media platform. This online “laboratory” was particularly useful because of its open automated programming interface (API); the free “Twitter Streaming API” provided researchers with up to 1% of all tweets, while the paid “Twitter Firehose” allowed researchers to extract 100% of all public posts [49]. As Twitter grew in size, the Firehose API was discontinued and replaced with the Decahose API which allowed researchers to download 10% of all posts in real time [12, 40]. The relative availability of Twitter data, along with the platform’s popularity and use across demographic and sociographic groups [47] spurred scientific innovations across a number of disciplines, including epidemic modeling [22], political communication [32], and statistical physics [24], to name just a few. Some sub-disciplines that depend on access to “everyday” communication, such as social movement studies, were revolutionized by access to the Twitter API, which provided data to support hundreds of papers on the structure and discursive content of social movements around the world (e.g., see Refs. [21, 29, 66]).
@@
The change in ownership of Twitter, subsequently known as X, spawned interest in many alternative platforms such as Mastodon, Threads, and Bluesky [39, 57, 59]. We focus our attention of Bluesky because of its inherent open access to platform data. Bluesky originated in 2019 as a research initiative at Twitter examining the potential for open-sourcing and decentralizing social media [54]. It was incorporated as an independent company in 2021  [45]. Bluesky’s central innovation is a standard known as the Authenticated Transfer (AT) Protocol [31]. The AT Protocol emphasizes federated identities, allowing users to own and curate a single online identity that is usable on all AT Protocol platforms and services [8]. User data is stored in repositories on personal data servers (PDSs) that log user actions; these repositories are intended to be portable so that users can easily migrate between hosts [9]. The Bluesky application queries a user’s PDS to deliver content to users; however, anyone can build an application to view AT Protocol data. Applications with access to the same data can vary in their content moderation policies, post visibility, or algorithmic curation strategies, for example [10].
@@
This Bluesky dataset addresses both of these gaps in the corpus of higher-order network datasets. In the summer of 2024, Bluesky developers announced a new feature in the Bluesky app, known as starter packs (SPs). Starter packs are user-curated lists of accounts and feeds (which serve post content to users based on custom algorithms) with a user-supplied title and description. Users share these lists so that other users can quickly grow the list of accounts they follow based on specified interest. One has the option to either browse the list and cherry-pick accounts to follow or hit the “follow all” button to follow all accounts in the list. Starter packs are inherently higher-order. At the time of this writing, users can create starter packs comprising 8 to 150 users and feeds. One caveat, however, is that a similar higher-order structure on Bluesky known as “lists”, which are not size restricted and so contain anywhere between zero and several thousand items. With the advent of starter packs, users developed tools for converting lists into starter packs, bypassing the standard size limits. So while the list of starter packs is almost entirely higher-order, there are exceptions to this (as can be later seen in Fig. 3). One can also extract the list of accounts that each user follows, which can be represented as a directed, pairwise network. These two types of social relationships bridge the gap between pairwise and higher-order network data. Furthermore, this dataset offers researchers ground to investigate social phenomena exclusive to groups and develop measures and algorithms tailored to such structures. In particular, the absence of edges in the following network when compared to the network induced by starter packs facilitates exploration of patterns that are exclusive to higher-order networks. For example, one can analyze user overlap patterns between starter packs, a structural feature inaccessible to pairwise network analysis. Higher-order network analysis enables a deeper understanding of the complex dynamics among users participating in group interactions.
Beyond applications for the study of higher-order dynamics and interactions, our data enables novel research in communication and human-computer interaction. Future work in this area can leverage our dataset to address questions of structural differences in starter pack inclusion, the effect of starter pack inclusion on following relationships, and more. This dataset provides an ideal testbed for developing machine learning algorithms which use pairwise features to learn higher-order network structure and vice-versa. Engineering systems to solve these tasks can unlock new insights into the relationship between peer-to-peer social networks and group-level interactions online. This dataset can also supplement the suite of higher-order social datasets used to validate group formation mechanisms and simulate the higher-order dynamics of information, norms, and social capital. Historical online social network datasets have been used to answer diverse research questions across a wide range of disciplines and our Bluesky dataset adds to this rich corpus, providing a unique opportunity to gain early insights into a key, distinguishing feature of a burgeoning new social media platform.
Methods
Collecting the data records was a multi-step process, requiring three steps for collecting the raw data from the Bluesky API as well as a final collation step. This process is illustrated in Fig. 1.
Figure 1: Illustration of the data collection pipeline. We export records from the Bluesky PLC Directory with the export endpoint to obtain the list of all possible PDSs and the list of possible DIDs with associated creation times. For each PDS in the list of PDSs obtained from the PLC Directory, we obtain the list of all DIDs hosted by that PDS along with their status with the listRepos endpoint. For active DIDs, we obtain the associated user repository containing all following and starter pack data using the getRepo endpoint. We then collate these three different records to obtain node, starter pack, and following datasets.
All user accounts on Bluesky have an associated decentralized identifier (DID). DIDs provide a persistent global reference for all account information and all queries to the Bluesky API to collect user information such as the accounts followed or the list of starter packs require a DID as a query parameter. DIDs allow us to link records in both the querying and collation steps.
Exporting records from the PLC Directory
The complete list of DIDs and all operations performed on these DIDs, such as creation, modification, and deletion, are stored in Bluesky’s “Public Ledger of Credentials (PLC)” hosted on the DID PLC Directory (https://plc.directory), a resource that is provided and maintained by Bluesky itself. A DID in the PLC directory, however, may not have an associated user account; as of Oct. 18, 2025, the PLC Directory contained 65,146,797 unique DIDs, in contrast to the 38,939,515 DIDs hosted on PDSs and the 39,650,447 accounts present in our dataset. However, the PLC Directory is useful for two reasons: first, the first appearance of a DID in the directory corresponds with its creation time and second, the directory lists all of the personal data servers (PDSs) on which user account data is hosted.
The export endpoint returns a cursor which allowed us to page through in batches of 1,000 records and download the contents of the PLC Directory, which took roughly 24 hours. We queried the PLC Directory for all entries with time stamps before October 18th, 2025 12:00:00 UTC. From this log file, we collected the list of possible PDS instances and the list of all possible DIDs with associated creation times, with the knowledge that they may not all be valid user accounts.
Querying the Personal Data Servers
While DIDs listed in the PLC Directory may not correspond to a Bluesky account, DIDs listed by a PDS are guaranteed to be active or inactive user accounts and we use this list of DIDs as our node list. Our method for querying user IDs by getting the list of DIDs from the PDS instances is more efficient for two reasons: first, we do not consider the 25M+ DIDs which have no associated user account, and second, querying the list of all accounts in a PDS along with their activity status allows us to avoid querying inactive accounts.
Querying the PDSs is a two-step process; first, the list of active user repositories hosted by a particular PDS is obtained and second, these user repositories are downloaded.
Obtaining the list of users in each Personal Data Server
While in principle one could collate the list of all DIDs by starting from a well-connected user account and performing a breadth-first search through the following relationships, the network is disconnected and, as seen in Fig. 6(c), some of the strongly-connected components are non-negligible in size. For this reason, we chose to first compile the list of accounts from the accounts listed across all PDSs and collect all following relationships and starter packs from each account.
We populated an asynchronous queue with PDS addresses and attempted to query each PDS in this queue through the com.atproto.sync.listRepos. Querying this endpoint returns a cursor which allowed us to page through in batches of 1000 DIDs and collect the entire list of accounts hosted by that PDS. Upon an unsuccessful query, we retried three times with exponential back-off if the status was 408, 429, 500, or 502-504. Of the 7,731 PDS addresses obtained from the PLC directory, only 3,163 were able to be queried, while 4,568 could not be queried. While we cannot be absolutely sure that these repositories cannot be queried, we manually double-checked a number of them. The output of the listRepos endpoint is in JSON format and the “repos” field contains a list of JSON objects, where each entry contains a DID, a boolean indicating whether the associated account is active, and, if not, the reason for its inactivity. Accounts become inactive either because the platform suspended their account or the user deactivated their account.
We obtain each DID’s creation time from the PLC directory as described above and if the DID was created prior to October 18th, 2025 12:00:00 UTC, we add the node to the list of nodes. This list of nodes contained the DID, the creation time, whether it was active, and its activity status (We added a “normal” status if it was active). If the DID was also active, we added the DID with its associated PDS address to another asynchronous queue which was used to query the user repositories. Interestingly, several PDSs contained identical lists of user accounts and in this case, we added the first occurrence to the queue and ignored all other occurrences.
Obtaining user records
We attempted to obtain the associated repository of each DID in the asynchronous queue of (DID, PDS) combinations generated in the previous step. For each active DID, we queried the com.atproto.sync.getRepo endpoint exposed by its host PDS. As with the listRepos endpoint, upon an unsuccessful query, we retried three times with exponential back-off if the status was 408, 429, 500, or 502-504. Of the 36,689,265 user repositories queried, 36,485,432 queries were successful and 203,833 queries were unsuccessful. Unsuccessful queries could be due to a number of factors, including permission issues which we noticed on a number of accounts, as well as server errors at the time of querying. Nonetheless, we note that the repositories unable to be queried comprise roughly 0.6% of the total, indicating a robust collection pipeline. This repository contains all user actions, such as posts, reposts, likes, blocks, follows, and list/starterpack creation. However, the repository is stored in Content Addressable aRchive (CAR) format, a binary format for efficient data storage and processing. We used a modified version of the atmst Python package to convert user records from this binary format to JSONL format.
We saved each user repository as a separate JSONL file, filtering on the creation time of the record; we only keep records created prior to October 18th, 2025 12:00:00 UTC. We queried the list of accounts for two weeks, ending on Nov 2, 2025.
Technical details
Because HTTP requests are an I/O limited operation, the Python asyncio and aiohttp packages offer an efficient way to query the Bluesky API. We used 1,024 asyncio tasks to consume both the PDS queue and the queue of (DID, PDS) combinations, and in accordance with industry wisdom, we limited the number of per-host concurrent connections to 64. We implemented a rate limiter which extracted the rate limits for each PDS from the response headers and ensured that the number of requests to each PDS was under this limit. We paired this rate limiter with retry logic, using an exponential back-off if the initial query was unsuccessful.
Collating records
As described in Fig. 1, we combine three separate data sources to create the node, following, and starter pack datasets.
Collating the starter pack network
Starter packs are stored in two types of user records: one describing starter pack-level metadata, such as the creator and creation time, and another listing the DIDs and the time at which they were added to the starter pack. We iterated through each user repository searching for app.bsky.graph.starterpack or app.bsky.graph.listitem records. From the former, we extract the creation date, the starter pack creator, and the list identifier associated with the starter pack members. From the latter, we collate lists of potential starter pack members comprising user accounts and feeds by their list identifier. We then pair list members with the starter packs by joining on the list identifier. Note that entities in a starter pack are not necessarily users, they can be any entity that can be “followed” on Bluesky (e.g. a list or a feed). Lists can only be followed, however, by circumventing the standard user interface through the API.
We parallelized the reading of each user repository across 64 processes and placed each starter pack stored as a Python dictionary in a queue to be consumed by a single-threaded file writer, since writes are not atomic on a distributed file system. The file writing operation saved the starter packs as a JavaScript Object Notation Lines (JSONL) file, where each starter pack is a line in the file. We used the XGI Python package [35] to convert the starter pack network to the HIF standard [16] and a hyperedge list.
Collating the following network
User following events are stored in app.bsky.graph.follow user records. Each record contains the following account, the account followed, and the time at which the follow occurred. We parallelized the reading of each user repository across 64 processes and placed the list of all follows as a comma-separated string where each line is a following event in a queue to be consumed by a single-threaded file writer. The file writing operation saved the following dataset as a Comma Separated Variables (CSV) File, where each following event is a line in the file.
Collating the account list
While in principle, the account list should be completely specified by the list of DIDs aggregated across all PDS instances, in practice, the following and starterpack datasets refer to DIDs not listed in the PDSs. The most plausible explanation is that we unsuccessfully queried the hosting PDS of this DID; despite retrying failed queries, it is possible that there are active servers that we were not able to query. We collate the list of all accounts by compiling lists of (1) all accounts listed in the following dataset, (2) all accounts listed in the starter pack dataset, and (3) all accounts listed by hosting PDSs and then taking the set union between these three lists. Because there are accounts in the following network and accounts in the starter pack network which are not listed by hosting PDSs, 791,650 accounts do not have statuses. For each DID in the collated list of DIDs, we start by attempting to pair this DID with a record from the listRepos record since this offers the richest data and if this record does not exist, we fall back to the creation time extracted from the PLC directory logs.
Anonymizing user data
@@ Methods
Although all user data is available via the public Bluesky API and users can require authentication to access their public user data from the API, we nonetheless take reasonable steps to prevent re-identification of users. We recognize, however, that with enough effort, that it is possible to re-identify all users. However, we assert that this effort will be comparable to querying the API from scratch. We anonymize our dataset in three ways: (1) by converting the DIDs of all users and feeds to unique integer identifiers, (2) by removing the name and description of all starter packs, and (3) by coarsening the precision of all timestamps to one day. We anonymize the accounts in the starter pack network, the following network, and the list of all accounts by mapping each DID to an integer. We do the same for each starter pack URI, converting it to an integer. Then we iterate through every starter pack and every following relationship using this mapping from DIDs and URIs to unique integers. This process ensures that while users cannot be identified, node IDs are consistent across the account list, the following network, and the starter pack network.
Limitations
@@
The approach we utilized in collecting the starter pack and following data comes with limitations, however. The number of active accounts, starter packs, and following relationships on the Bluesky platform is rapidly changing and any collection method not taking this into account will not perfectly capture the network structure. Using the PLC Directory as a static directory does not consider accounts created after this list was collected. This could potentially lead to missing accounts that have been migrated to a PDS created after the last entry of the PLC directory collected, but before the accounts were successfully queried. Because we remove all events with creation dates after October 18th, 2025 12:00:00 UTC, we will not, in principle, collect network data for DIDs absent from the PLC directory; however, we are unable to collect accounts that are deactivated, starter packs that are deleted, and accounts that are unfollowed before they have been successfully queried. Despite this temporal filtering, however, there are nonetheless a small number of accounts and events with timestamps later than this cutoff time. We suspect that this could be a server-side issue or a problem with time zone conversion. However, they are also associated with following events which fall within the specified cutoff time. To this point, time stamps and user data can be arbitrarily set by users if they interact with Bluesky directly through the AT Protocol and not through the Bluesky application. We discuss this limitation in the Technical Validation section. It is also not strictly true that the network data we collect is in the PLC directory; there are 3,441 accounts in the list of accounts we collate which are not present in the PLC Directory. One way around some of these limitations is by downloading streaming data from the Bluesky Firehose API while backfilling all events for all users, as is done in Ref. [5]. The downside to this approach is that it is an immense amount of data; roughly 27GB of JSON Lines (JSONL) data per day at the time of writing. We believe that our approach strikes the balance of being computationally manageable, while surveying close to the entire network.
As the Bluesky social media platform continues to grow and evolve, it is important to augment this dataset with new data. Our method can be used to periodically create “snapshots” of the network, but this raises the important question of the timescales of these dumps. Ref. [5] noted that over the course of their study, 20% of starter packs were deleted. Our static method cannot observe these deletions, nor the precise time at which they occur. Long term, the most robust method for complete temporal data is through a pipeline using the Bluesky Firehose API, but it is beyond the scope of this dataset.
Ethical Considerations
@@
This data collection was ruled an exempt protocol by the Institutional Review Board (IRB) at Northeastern University (No. 25-01-22) and by the Institutional Review Board for the Social and Behavioral Sciences (IRB-SBS) at the University of Virginia (No. 7332). All of the data obtained is publicly accessible (without authentication or a Bluesky account) via the Bluesky API. Additionally, this study was granted a waiver of informed consent from both the IRB at Northeastern University and the IRB-SBS at the University of Virginia; arguably, obtaining informed consent for all users’ inclusion in this dataset would constitute a larger risk to users’ privacy than the collection of publicly available data. Nonetheless, we recognize that this dataset could enable large-scale identification of accounts’ following relationships. We therefore redact all user DIDs (decentralized identifiers, used for long-term, persistent identification of entities in the AT Protocol that Bluesky uses) and truncate the timestamps on all starter pack and account creation events to include only the date; i.e., the level of precision at which timestamps are reported is one day. The details of this process are described in the section detailing the anonymization of user data.
Data Records
This study releases a list of anonymized accounts and two large-scale networks: (1) a list of starter packs on Bluesky and (2) a list of user following relationships. All data is available on the Social Media Archive (SOMAR), which is a project by the Inter-university Consortium for Political and Social Research. The datasets described in this paper are available in the data collection titled “A Blue Start: A large-scale pairwise and higher-order social network dataset” [62]. The dataset is available under a CC-BY license with an additional clause that stipulates users will not attempt to reidentify individuals in the dataset.
List of nodes
We provide a list of nodes in the network with accompanying metadata for each node.
JSON Lines (deidentified_nodes.jsonl.gz, 221.4 MB). Each line in this JSON Lines file is a JSON dictionary with mandatory fields “node-id,” representing the node’s anonymized integer identifier, and “active,” which is a Boolean attribute indicating whether the node was active or taken down at the time of data collection. The optional fields provide further information about the nodes: “date-created” is the date (in YYYY-mm-dd format) on which the node was created; status is one of “deactivated,” “normal,” or “takendown,” indicating whether the node is inactive (initiated by the user), active, or inactive (initiated by the platform), respectively.
CSV (deidentified_nodes.csv.gz, 187.7 MB). Each line in this compressed CSV file contains the fields “id,” “date-created,” “active,” and “status.” The “id” field represents the node’s anonymized integer identifier; the “date-created” field is the date (in YYYY-mm-dd format) on which the node was created; the “active” field is a Boolean attribute indicating whether the node was active or taken down at the time of data collection; and the “status” field is one of “deactivated,” “normal,” or “takendown,” indicating whether the node is inactive (initiated by the user), active, or inactive (initiated by the platform), respectively.
Bluesky starter packs
@@
JSON Lines (deidentified_starterpacks.jsonl.gz, 86.1 MB). We provide a JSON Lines file of all scraped starter packs’ contents and metadata, intended for computational analysis. This file contains a list of starter packs; each starter pack contains the fields described in Table 1.
@@
members array List of starter pack members
Entries in the members array
id integer The member’s anonymous integer identifier
date-added date The date on which the member account was added to the starter pack
@@
Hypergraph Interchange Format (deidentified_starterpack_hif.json.gz, 111.5 MB). This file is a HIF-compliant JSON file. The HIF standard provides a cross-platform standard for sharing attributed higher-order network data. The edge attributes (“id”, “creator-id”, and “date-created”) are labeled as described in Table 1. Each entry in the “incidences” field has attributes “edge,” corresponding to a starter pack (edge) ID, “node,” corresponding to a node ID, and “date-added,” indicating when the node was added to the starter pack, as explained in Table 1. The attributes included in the “nodes” field are “node” (node ID), “active,” and “status” as seen in the node list described previously.
Hyperedge list (deidentified_starterpack_edgelist.csv.gz, 46.6 MB). We also provide a hyperedge list stored as a gzipped CSV file. Each line in this file corresponds to a starter pack, and each entry in the line is a member of the starter pack. Each line contains entries separated by commas; a single line corresponds to a single starter pack. For example, the line “1,2,3,4” indicates that users with anonymous integer identifiers 1, 2, 3, and 4 belong to the same starter pack.
Bluesky following network
We also release the following network in two different formats.
Edgelist CSV (deidentified_follows_edgelist.csv.gz, 14.9 GB). We provide the following network dataset as a compressed CSV file. Each line $(i,j,d)$ in the file indicates that user $i$ followed user $j$ on date $d$, where $d$ is provided in YYYY-mm-dd format. Here, $i$ and $j$ are anonymous integer identifiers that map to user DIDs. As described in the section detailing the anonymization of user data, these anonymous integer IDs can be linked to creators and members of starter packs.
Edgelist Parquet (deidentified_follows_edgelist.parquet, 12.6 GB). This data artifact has the same format as the CSV, but it is encoded in the more efficient Parquet format and has column names from, to, and date_followed.
Technical Validation